My age is thrice as my brother age. The difference of our age is 18. What is my age?
Answer by advanced_Learner(501)

My age is thrice as my brother age. The difference of our age is 18. What is my age?
let x be my brother age and y be my age
y=3x
y-x=18
3x-x=18
2x=18
x=9 brother age
y=3*9
y=27 is my age
enjoy